OpenStudy (anonymous):

Using a trig identity, write y(t)= -2cos(4t)-sin(4t) using only one cosine function

OpenStudy (anonymous):



OpenStudy (anonymous):

Use auxilliary angles to write it in the form of \[\large R\cos(\epsilon+4t)\] where \[\large \epsilon is the angle you are finding. \[\large R=\sqrt{(-2)^2+(-1)^2}\] \[\large R=\sqrt{5}\] \[\large \sqrt{5}[-\frac{2}{\sqrt{5}}\cos(4t)-\frac{1}{\sqrt{5}}\sin(4t)]\] Therefore \[\large \cos (\epsilon) =-\frac{2}{\sqrt{5}}\] \[\large \epsilon=153`26'\] NOTE:` denotes a degree sign. I don't have one on my keyboard. THEREFORE: \[\huge -2\cos(4t)-\sin(4t)= \sqrt{5}\cos(153`26'+4t)\]

OpenStudy (anonymous):

\[\huge y(t)=\sqrt{5}\cos(153`26'+4t)\]

OpenStudy (anonymous):

\[put r \cos \theta=-2,r \sin \theta=1\] square and add \[r ^{2}\left( \cos ^{2}\theta+\sin ^{2} \theta\right)=4+1=5\] \[r ^{2}=5,r=\sqrt{5}\] divide, \[\tan \theta=-\frac{ 1 }{ 2 },\theta=\tan^{-1} \left( \frac{ -1 }{ 2 } \right)\] \[y \left( t \right)=r \cos \theta \cos \left( 4t \right)-r \sin \theta \sin \left( 4t \right)=r \cos \left( 4t+theta \right)\]

OpenStudy (anonymous):

put the values of r & theta ,we get the result

OpenStudy (anonymous):

y"+y'=2 \[\frac{ d }{ dx }\left( \frac{ dy }{dx } \right)+\frac{ dy }{dx }=2\] integrating w.r.t. x, \[\frac{ dy }{dx }+y=2x+c\] it is an exact diff. eq. \[I.F=e ^{\int\limits1 dx}=e ^{x}\] \[c.s. is y*e ^{x}=\int\limits \left( 2x+c \right)e ^{x}dx+c1\] \[y e ^{x}=c \int\limits e ^{x} dx+2\int\limits x e ^{x} dx+c1\] \[y e ^{x}=c e ^{x}+2\int\limits x e ^{x} dx+c1\] \[\int\limits xe ^{x}dx=first function *\left( \int\limits of second function \right) \[-\int\limits \frac{ d }{ dx } first function*\int\limits second function dx dx\] \] \[\int\limits x e ^{x} dx=x \int\limits e ^{x}dx-1\int\limits e ^{x}dx=x e ^{x}-e ^{x}=\left( x-1 \right)e ^{x}\] substitute and get the complete solution.